# ExaModelsPower.jl: A GPU-Compatible Modeling Library for Nonlinear Power System Optimization

### v1

@@ I Introduction
Alternating current (AC) optimal power flow (OPF) is used to determine optimal power system operating points, which must satisfy nonlinear power flow constraints and various operational constraints [1, 2, 3]. Resolving AC OPF problems, formulated as nonlinear programming (NLP), can be computationally challenging, especially for large networks with thousands of buses, lines, and generators. The development of efficient solution methods for these problems remains an active area of research. In practice, system operators (SO) often solve direct current (DC) OPF problems, i.e., linear approximations that neglect the inherent nonlinearities of the power flow equations [1, 4]. Such approximations can lead to suboptimal or infeasible solutions as DC feasibility does not guarantee AC feasibility. It is estimated that switching from DC to AC OPF could reduce electricity costs by several percentage points, translating into billions of dollars in yearly savings for the US [1]. Numerous efforts, such as the ARPA-E Grid Optimization Competitions [5, 6, 7, 8], have thus been undertaken to improve solution methods for AC OPF. Results from the competition show that solving the nonlinear power flow equations remains a computational bottleneck for large-scale systems [9, 10, 11].
@@
Capacity expansion planning determines investments in generation, transmission, and storage assets over multi-year horizons. Software tools for capacity expansion planning include GenX [17], ReEDS [18], and PyPSA [19]. Unit commitment determines the scheduling of generation units, which typically involves logical constraints modeled by auxiliary binary variables. Software tools for unit commitment include PLEXOS [20] and UnitCommitment [21].
AC OPF determines power system operating points that respect all operational constraints, including the nonlinear power flow equations, at least cost. PowerModels.jl and MATPOWER provide diverse options for user-specified AC OPF models, respectively in Julia and Matlab [3, 15]. PowerModels.jl uses JuMP [22] as a modeling backend, while MATPOWER implements its own custom derivative evaluation. PYPOWER ports MATPOWER to allow for OPF modeling in Python [23]. Pandapower is built on pandas—a data analysis library in Python—and provides a Python-based tool for power system modeling and analysis [24]. Pandapower can be used for optimization, but relies on PowerModels.jl or PYPOWER as external modeling backends.
ExaGo provides the most similar functionality to ExaModelsPower.jl, as it also aims to provide an interface for creating GPU-compatible AC OPF models [25]. ExaGo uses PETSc [26] as modeling backend with RAJA [27] and Umpire [28] for GPU memory management and loop abstractions. To our knowledge, in their current version v1.6.0, AC OPF models cannot be solved in a fully GPU-resident manner because none of the solvers interfaced to ExaGo are fully GPU-compatible. For example, HiOp currently handles mixed sparse-dense formulations and does not support the fully sparse formulations required for large-scale AC OPF problems [29]. We are not aware of any published performance benchmarking of ExaGo running on GPU.
@@ II-D Modeling Syntax
Similarly to LABEL:model_exa, the model is constructed by first defining the variables, objective, and constraints using the ExaModels.jl standard API. Note that constraint! is used to modify an existing constraint, which is useful when the constraint needs to be modified multiple times, such as in the case of power balance constraints where the contributions from various devices are summed up. An ExaModel object is then created using the core object that contains all of the cached model information. ExaModelsPower.jl implements all the internal functions required to model the OPF in both polar and rectangular coordinates, such as gen_cost, c_ref_angle_polar, c_to_active_power_flow_polar, etc.
@@ III-C1 Static OPF Models
model, vars, cons = opf_model("pglib_opf_case118_ieee.m";)
@@
model, vars, cons = opf_model("pglib_opf_case118_ieee.m"; backend = CUDABackend(), form = :rect)
@@ III-C4 GOC3 Models
model, vars, cons = goc3_model("data/C3E4N00073D1_scenario_303.json", "data/C3E4N00073D1_scenario_303_solution.json")
@@ IV-A3 Solver Configurations
  • Ipopt (C)—Ipopt v3.14.19 with ma27 on CPU [40]: Ipopt is a widely used open-source CPU-based interior-point solver. When it is configured with HSL ma27 linear solver, the entire solver runs on a single thread. For static AC OPF, Ipopt with ma27 is among the fastest open-source solvers [47]. Ipopt can reliably find the local optimal solution of static AC OPF problems with $10^{-8}$ tolerance (based on relative error to KKT conditions).
  • MadNLP (LK, C)—MadNLP.jl v0.8.10 with LiftedKKT and ma86 on CPU [14]: MadNLP.jl is an interior-point solver that can run on both CPU and GPU, and its algorithm is based on the filter line-search method used by Ipopt [40]. MadNLP.jl supports various KKT systems beyond the augmented system used in Ipopt, including the LiftedKKT [14] and HybridKKT [36]. LiftedKKT is particularly well-suited for linear solvers that exploit multi-thread parallelism at the elimination-tree level, such as ma86, because the matrix treated by direct solver is always made positive definite, thus eliminating the need for numerical pivoting [14]. In this configuration, MadNLP.jl runs on a single thread, but the linear solver ma86 can utilize multiple threads. LiftedKKT slightly relaxes the equality constraints by the solver tolerance; thus, when loose tolerance is used, the final solution may have a non-negligible constraint violation.
  • MadNLP (LK, G)—MadNLP.jl v0.8.10 with LiftedKKT and cuDSS on GPU [14]: This configuration runs on an NVIDIA GPU and solves sparse linear systems using cuDSS. By default, MadNLP.jl uses the LiftedKKT formulation when the backend is set to GPU. MadNLP.jl shares a common code base for CPU and GPU versions. However, the low-level kernels are specialized for the hardware backend via Julia’s multiple dispatch, ensuring that the performance-critical parts are optimized for GPU execution. Thus, this configuration of solver will run the same algorithm as the MadNLP+LiftedKKT+ma86 CPU configuration, but everything is executed on the GPU instead.
  • MadNLP (HK, G)—MadNLP.jl v0.8.10 with HybridKKT and cuDSS on GPU [14]: HybridKKT [36] is used in this configuration instead of LiftedKKT. Most algorithmic details are the same as in the LiftedKKT GPU configuration. HybridKKT does not apply the equality relaxation, and thus, is able to achieve better constraint satisfaction than LiftedKKT for the same convergence tolerance; however, HybridKKT internally relies on an iterative solver to solve the Schur complement system, which may lead to slower performance than LiftedKKT in some cases.
  • MadNCL (G)—MadNCL.jl v0.1.0 with cuDSS on GPU [48]: MadNCL.jl is a new augmented-Lagrangian–based solver that runs on both CPU and GPU. Like MadNLP.jl, MadNCL.jl shares a common code base for CPU and GPU versions, but low-level kernels are compiled differently for each hardware backend.
@@ IV-B Static OPF Formulation
Table I presents the solver performances for the static OPF modeled in polar and rectangular coordinates. For both setups, we use a 900-second wall-clock time limit. Figure 1 shows the speedup achieved by different solvers as problem size grows, compared to Ipopt configured with ma27, running on CPU.
Key observations are as follows: (1) When the SGM10 time is compared across all problem sizes, MadNLP+LiftedKKT yields the best performance for moderate tolerance, but MadNCL performs better when a polar formulation is used and a tight tolerance of $10^{-8}$ is required. (2) When we focus on large problems only, MadNLP on GPU is the fastest for both polar and rectangular formulations, and for both tolerances. (3) For small problems, CPU solvers are consistently faster than GPU solvers. (4) All three GPU solvers experience more frequent failures under tighter tolerances, while CPU solvers show only one additional failure. (5) Solve times are almost always shorter when the problem is modeled in polar coordinates, except for small problems.
@@
Polar
$10^{-4}$ MadNLP (LK, G) 23 0.82 1.38e-3 31 1.47 1.47e-3 12 3.42 2.18e-3 66 1.57 1.57e-3
MadNLP (HK, G) 22 2.49 8.76e-7 23 24.0 2.61e-7 12 4.45 1.21e-7 57 10.53 4.69e-7
MadNCL (G) 21 5.64 8.70e-5 31 2.32 5.38e-3 12 10.73 8.62e-4 64 4.72 2.80e-3
Ipopt (C) 23 0.25 1.88e-6 31 3.75 7.57e-6 12 33.26 5.37e-6 66 5.29 5.18e-6
MadNLP (LK, C) 22 2.46 1.93e-6 31 10.72 1.05e-6 11 50.16 8.64e-7 64 11.06 1.32e-6
$10^{-8}$ MadNLP (LK, G) 22 2.86 1.36e-7 31 4.12 1.45e-7 9 35.25 1.17e-7 62 6.89 1.38e-7
MadNLP (HK, G) 22 2.53 1.35e-7 23 25.22 1.28e-7 10 22.22 2.41e-7 55 14.17 1.51e-7
MadNCL (G) 19 13.34 1.55e-7 31 3.0 2.06e-7 7 87.63 1.16e-7 57 13.0 1.78e-7
Ipopt (C) 23 0.27 1.35e-7 31 4.26 1.44e-7 12 37.23 2.18e-7 66 5.81 1.55e-7
MadNLP (LK, C) 22 2.66 1.37e-7 31 15.82 1.46e-7 10 88.27 1.20e-7 63 15.68 1.38e-7
Rectangular
$10^{-4}$ MadNLP (LK, G) 23 0.4 1.39e-3 30 3.09 1.46e-3 10 18.47 2.36e-3 63 3.91 1.58e-3
MadNLP (HK, G) 22 2.55 9.89e-7 22 29.51 1.39e-6 9 31.64 1.43e-7 53 16.75 1.01e-6
MadNCL (G) 21 5.43 1.08e-4 29 6.86 3.92e-3 5 163.18 1.89e-3 55 14.97 2.28e-3
Ipopt (C) 23 0.17 1.98e-6 30 6.29 8.83e-6 9 78.23 4.95e-6 62 8.79 5.73e-6
MadNLP (LK, C) 23 0.86 1.64e-6 30 12.1 1.26e-6 8 115.17 1.08e-6 61 13.65 1.38e-6
$10^{-8}$ MadNLP (LK, G) 22 2.73 1.37e-7 29 6.36 1.40e-7 8 50.44 2.26e-7 59 9.01 1.51e-7
MadNLP (HK, G) 22 2.62 1.35e-7 22 31.62 1.28e-7 7 86.01 2.85e-7 51 21.97 1.53e-7
MadNCL (G) 21 5.64 1.31e-7 29 7.56 1.60e-7 4 236.94 1.06e-7 54 17.27 1.45e-7
Ipopt (C) 23 0.2 1.35e-7 30 6.76 1.45e-7 9 83.34 2.47e-7 62 9.26 1.56e-7
MadNLP (LK, C) 23 1.21 1.38e-7 30 14.14 1.45e-7 8 134.84 2.26e-7 61 15.59 1.53e-7
@@ IV-C MP OPF Formulation (with or without storage)
Table II presents the benchmark results for the MPOPF cases with and without storage. Figure 2 depicts the speedup achieved by different solvers as problem size increases for the MPOPF without storage. The results for the MP OPF benchmarking are summarized as so: (1) Due to the time limit, GPU solvers have much higher success rate than CPU solvers, particularly for large instances. Furthermore, within the given time limit, large-sized cases with storage are only solvable using GPU solvers. (2) Adding storage decreases the amount of cases that any solver can solve, especially for large cases at high tolerance. A large portion of these failures arise from max wall time violations. (3) GPU solvers achieve nearly 2 orders of magnitude speedup for the largest instances that are solvable on CPU in the given time limit. (4) At both tolerances, multi-thread CPU parallelism via ma86 provides substantial speed-ups over the single-threaded solver Ipopt with ma27. (5) LiftedKKT and MadNCL.jl exhibit larger constraint violations compared to the other solvers.
@@
No Storage
$10^{-4}$ MadNLP (LK, G) 13 0.4 5.28e-4 9 7.09 2.59e-3 18 59.41 1.95e-3 40 17.33 1.63e-3
MadNLP (HK, G) 13 0.62 8.19e-7 8 18.09 7.34e-7 12 215.9 7.57e-7 33 42.32 7.76e-7
MadNCL (G) 13 0.63 2.36e-5 8 51.22 1.14e-2 14 225.28 3.89e-3 35 53.51 4.16e-3
Ipopt (C) 13 0.66 2.49e-6 7 132.77 8.41e-6 1 894.61 1.03e-6 21 131.0 4.40e-6
MadNLP (LK, C) 13 1.53 1.07e-6 8 86.38 4.93e-7 2 832.88 1.40e-5 23 118.27 1.99e-6
$10^{-8}$ MadNLP (LK, G) 13 0.95 5.28e-8 8 13.64 2.70e-7 12 239.19 1.56e-7 33 35.57 1.43e-7
MadNLP (HK, G) 13 0.52 4.99e-8 6 40.03 1.83e-7 8 259.11 1.86e-7 27 44.95 1.20e-7
MadNCL (G) 12 5.0 1.74e-7 5 89.39 4.15e-7 5 476.13 5.53e-7 22 82.91 3.15e-7
Ipopt (C) 13 0.89 4.95e-8 6 126.28 3.08e-7 0 - - 19 103.93 1.31e-7
MadNLP (LK, C) 13 3.39 5.28e-8 7 96.23 2.94e-7 1 833.42 4.19e-7 21 102.73 1.51e-7
Storage
$10^{-4}$ MadNLP (LK, G) 16 3.13 5.70e-4 8 100.59 1.15e-3 15 305.46 8.94e-4 39 60.65 8.14e-4
MadNLP (HK, G) 10 49.8 2.35e-6 0 - - 0 - - 10 304.52 2.35e-6
MadNCL (G) 16 2.35 1.26e-4 9 137.53 8.66e-3 8 714.84 5.88e-4 33 90.32 2.56e-3
Ipopt (C) 15 12.14 1.83e-6 4 627.02 9.81e-8 0 - - 19 185.63 1.46e-6
MadNLP (LK, C) 15 14.64 8.27e-7 5 387.67 1.78e-7 0 - - 20 171.85 6.65e-7
$10^{-8}$ MadNLP (LK, G) 14 11.87 5.51e-8 6 119.75 1.53e-7 1 861.23 8.84e-8 21 38.34 8.47e-8
MadNLP (HK, G) 12 28.26 2.72e-8 0 - - 0 - - 12 109.73 2.72e-8
MadNCL (G) 14 10.57 5.44e-8 4 297.69 1.14e-7 1 821.55 6.39e-8 19 48.98 6.73e-8
Ipopt (C) 15 16.68 5.14e-8 1 642.34 6.23e-8 0 - - 16 76.59 5.21e-8
MadNLP (LK, C) 16 10.1 5.66e-8 4 446.25 8.86e-8 0 - - 20 55.37 6.30e-8

### v2

@@ I Introduction
Alternating current (AC) optimal power flow (OPF) is used to determine optimal power system operating points, which must satisfy nonlinear power flow constraints and various operational constraints [1, 2, 3]. Resolving AC OPF problems, formulated as nonlinear programming (NLP), can be computationally challenging, especially for large networks with thousands of buses, lines, and generators. The development of efficient solution methods for these problems remains an active area of research. In practice, system operators (SO) often solve direct current (DC) OPF problems, i.e., linear approximations that neglect the inherent nonlinearities of the power flow equations [1, 4]. Such approximations can lead to suboptimal or infeasible solutions as DC feasibility does not guarantee AC feasibility. It is estimated that switching from DC to AC OPF could reduce electricity costs by several percentage points, translating into billions of dollars in yearly savings for the US [1]. Modeling withAC OPF not only leads to efficiency improvements but also improves reliability by considering the roles of reactive power and voltage magnitude. Numerous efforts, such as the ARPA-E Grid Optimization Competitions [5, 6, 7, 8], have thus been undertaken to improve solution methods for AC OPF. Results from the competition show that solving the nonlinear power flow equations remains a computational bottleneck for large-scale systems [9, 10, 11].
@@
Capacity expansion planning determines investments in generation, transmission, and storage assets over multi-year horizons. Software tools for capacity expansion planning include GenX [17], ReEDS [18], and PyPSA [19].
Unit commitment determines the scheduling of generation units, which typically involves logical constraints modeled by auxiliary binary variables. Software tools for unit commitment include PLEXOS [20] and UnitCommitment [21].
AC OPF determines power system operating points that respect all operational constraints, including the nonlinear power flow equations, at least cost. PowerModels.jl and MATPOWER provide diverse options for user-specified AC OPF models, respectively in Julia and Matlab [3, 15]. PowerModels.jl uses JuMP [22] as a modeling backend, while MATPOWER implements its own custom derivative evaluation. PYPOWER ports MATPOWER to allow for OPF modeling in Python [23]. Pandapower is built on pandas—a data analysis library in Python—and provides a Python-based tool for power system modeling and analysis [24]. Pandapower can be used for optimization, but relies on PowerModels.jl or PYPOWER as external modeling backends. ExaGo provides the most similar functionality to ExaModelsPower.jl, as it also aims to provide an interface for creating GPU-compatible AC OPF models [25]. ExaGo uses PETSc [26] as modeling backend with RAJA [27] and Umpire [28] for GPU memory management and loop abstractions. To our knowledge, in their current version v1.6.0, AC OPF models cannot be solved in a fully GPU-resident manner because none of the solvers interfaced to ExaGo are fully GPU-compatible. For example, HiOp currently handles mixed sparse-dense formulations and does not support the fully sparse formulations required for large-scale AC OPF problems [29]. We are not aware of any published performance benchmarking of ExaGo running on GPU.
@@ II-D Modeling Syntax
Similarly to LABEL:model_exa, the model is constructed by first defining the variables, objective, and constraints using the ExaModels.jl standard API. Note that constraint! is used to modify an existing constraint, which is useful when the constraint needs to be modified multiple times, such as in the case of power balance constraints where the contributions from various devices are summed up. For example, the syntax a.bus=>p[a.i] adds the term p[a.i] to the a.bus-th constraint in the original set of constraints. An ExaModel object is then created using the core object that contains all of the cached model information. ExaModelsPower.jl implements all the internal functions required to model the OPF in both polar and rectangular coordinates, such as gen_cost, c_ref_angle_polar, c_to_active_power_flow_polar, etc.
@@ III-C1 Static OPF Models
model, vars, cons = ac_opf_model("pglib_opf_case118_ieee.m";)
@@
model, vars, cons = ac_opf_model("pglib_opf_case118_ieee.m"; backend = CUDABackend(), form = :rect)
@@ III-C4 GOC3 Models
model, cons, vars, lengths, sc_data_array = goc3_model("data/C3E4N00073D1_scenario_303.json", "data/C3E4N00073D1_scenario_303_solution.json")
@@ IV-A3 Solver Configurations
  • Ipopt (CPU)—Ipopt v3.14.19 with ma27 on CPU [40]: Ipopt is a widely used open-source CPU-based interior-point solver. When it is configured with HSL ma27 linear solver, the entire solver runs on a single thread. For static AC OPF, Ipopt with ma27 is among the fastest open-source solvers [47]. Ipopt can reliably find the local optimal solution of static AC OPF problems with $10^{-8}$ tolerance (based on relative error to KKT conditions).
  • MadNLP (LK, CPU)—MadNLP.jl v0.8.10 with LiftedKKT and ma86 on CPU [14]: MadNLP.jl is an interior-point solver that can run on both CPU and GPU, and its algorithm is based on the filter line-search method used by Ipopt [40]. MadNLP.jl supports various KKT systems beyond the augmented system used in Ipopt, including the LiftedKKT [48] and HybridKKT [36]. LiftedKKT is particularly well-suited for linear solvers that exploit multi-thread parallelism at the elimination-tree level, such as ma86, because the matrix treated by direct solver is always made positive definite, thus eliminating the need for numerical pivoting. More details on the algorithmic differences between LiftedKKT, HybridKKT and the augmented system can be found in [48]. In this configuration, MadNLP.jl runs on a single thread, but the linear solver ma86 can utilize multiple threads. LiftedKKT slightly relaxes the equality constraints by the solver tolerance; thus, when loose tolerance is used, the final solution may have a non-negligible constraint violation.
  • MadNLP (LK, GPU)—MadNLP.jl v0.8.10 with LiftedKKT and cuDSS on GPU [14]: This configuration runs on an NVIDIA GPU and solves sparse linear systems using cuDSS. By default, MadNLP.jl uses the LiftedKKT formulation when the backend is set to GPU. MadNLP.jl shares a common code base for CPU and GPU versions. However, the low-level kernels are specialized for the hardware backend via Julia’s multiple dispatch, ensuring that the performance-critical parts are optimized for GPU execution. Thus, this configuration of solver will run the same algorithm as the MadNLP+LiftedKKT+ma86 CPU configuration, but everything is executed on the GPU instead.
  • MadNLP (HK, GPU)—MadNLP.jl v0.8.10 with HybridKKT and cuDSS on GPU [14]: HybridKKT [36] is used in this configuration instead of LiftedKKT. Most algorithmic details are the same as in the LiftedKKT GPU configuration. HybridKKT does not apply the equality relaxation, and thus, is able to achieve better constraint satisfaction than LiftedKKT for the same convergence tolerance; however, HybridKKT internally relies on an iterative solver to solve the Schur complement system, which may lead to slower performance than LiftedKKT in some cases.
  • MadNCL (GPU)—MadNCL.jl v0.1.0 with cuDSS on GPU [49]: MadNCL.jl is a new augmented-Lagrangian–based solver that runs on both CPU and GPU. Like MadNLP.jl, MadNCL.jl shares a common code base for CPU and GPU versions, but low-level kernels are compiled differently for each hardware backend.
@@ IV-B Static OPF Formulation
Table I presents the solver performances for the static OPF modeled in polar and rectangular coordinates. For both setups, we use a 900-second wall-clock time limit. Figure 1 shows the speedup achieved by different solvers as problem size grows, using the case of Ipopt configured with ma27 running on CPU as a baseline.
Key observations are as follows: (1) When the SGM10 time is compared across all problem sizes, MadNLP+LiftedKKT yields the best performance for moderate tolerance, but MadNCL performs better when a polar formulation is used and a tight tolerance of $10^{-8}$ is required. (2) When we focus on large problems only, MadNLP on GPU is the fastest for both polar and rectangular formulations, and for both tolerances. (3) For small problems, CPU solvers are consistently faster than GPU solvers. This result is expected, as the advantages of GPU computation are lower on small cases which allow for a smaller degree of parallelization. (4) All three GPU solvers experience more frequent failures under tighter tolerances, while CPU solvers show only one additional failure. (5) Solve times are almost always shorter when the problem is modeled in polar coordinates, except for small problems.
@@
Polar $10^{-4}$ MadNLP (LK, GPU) 23 0.82 1.38e-3 31 1.47 1.47e-3 12 3.42 2.18e-3 66 1.57 1.57e-3
MadNLP (HK, GPU) 22 2.49 8.76e-7 23 24.0 2.61e-7 12 4.45 1.21e-7 57 10.53 4.69e-7
MadNCL (GPU) 21 5.64 8.70e-5 31 2.32 5.38e-3 12 10.73 8.62e-4 64 4.72 2.80e-3
Ipopt (CPU) 23 0.25 1.88e-6 31 3.75 7.57e-6 12 33.26 5.37e-6 66 5.29 5.18e-6
MadNLP (LK, CPU) 22 2.46 1.93e-6 31 10.72 1.05e-6 11 50.16 8.64e-7 64 11.06 1.32e-6
$10^{-8}$ MadNLP (LK, GPU) 22 2.86 1.36e-7 31 4.12 1.45e-7 9 35.25 1.17e-7 62 6.89 1.38e-7
MadNLP (HK, GPU) 22 2.53 1.35e-7 23 25.22 1.28e-7 10 22.22 2.41e-7 55 14.17 1.51e-7
MadNCL (GPU) 19 13.34 1.55e-7 31 3.0 2.06e-7 7 87.63 1.16e-7 57 13.0 1.78e-7
Ipopt (CPU) 23 0.27 1.35e-7 31 4.26 1.44e-7 12 37.23 2.18e-7 66 5.81 1.55e-7
MadNLP (LK, CPU) 22 2.66 1.37e-7 31 15.82 1.46e-7 10 88.27 1.20e-7 63 15.68 1.38e-7
Rectangular $10^{-4}$ MadNLP (LK, GPU) 23 0.4 1.39e-3 30 3.09 1.46e-3 10 18.47 2.36e-3 63 3.91 1.58e-3
MadNLP (HK, GPU) 22 2.55 9.89e-7 22 29.51 1.39e-6 9 31.64 1.43e-7 53 16.75 1.01e-6
MadNCL (GPU) 21 5.43 1.08e-4 29 6.86 3.92e-3 5 163.18 1.89e-3 55 14.97 2.28e-3
Ipopt (CPU) 23 0.17 1.98e-6 30 6.29 8.83e-6 9 78.23 4.95e-6 62 8.79 5.73e-6
MadNLP (LK, CPU) 23 0.86 1.64e-6 30 12.1 1.26e-6 8 115.17 1.08e-6 61 13.65 1.38e-6
$10^{-8}$ MadNLP (LK, GPU) 22 2.73 1.37e-7 29 6.36 1.40e-7 8 50.44 2.26e-7 59 9.01 1.51e-7
MadNLP (HK, GPU) 22 2.62 1.35e-7 22 31.62 1.28e-7 7 86.01 2.85e-7 51 21.97 1.53e-7
MadNCL (GPU) 21 5.64 1.31e-7 29 7.56 1.60e-7 4 236.94 1.06e-7 54 17.27 1.45e-7
Ipopt (CPU) 23 0.2 1.35e-7 30 6.76 1.45e-7 9 83.34 2.47e-7 62 9.26 1.56e-7
MadNLP (LK, CPU) 23 1.21 1.38e-7 30 14.14 1.45e-7 8 134.84 2.26e-7 61 15.59 1.53e-7
Cvio = maximum constraint violation, nvar = Number of Variables, LK = LiftedKKT, HK = HybridKKT.
@@ IV-C MP OPF Formulation (with or without storage)
Table II presents the benchmark results for the MPOPF cases with and without storage. Figure 2 depicts the speedup achieved by different solvers as problem size increases for the MPOPF without storage. Figure 3 presents the speed of various solvers as a factor of the fastest solver across cases. The results for the MP OPF benchmarking are summarized as so: (1) Due to the time limit, GPU solvers have much higher success rate than CPU solvers, particularly for large instances. Furthermore, within the given time limit, large-sized cases with storage are only solvable using GPU solvers. (2) Adding storage decreases the amount of cases that any solver can solve, especially for large cases at high tolerance. A large portion of these failures arise from max wall time violations. (3) GPU solvers achieve nearly 2 orders of magnitude speedup for the largest instances that are solvable on CPU in the given time limit. (4) GPU solvers generate the fastest solution for approximately two-thirds of cases. (5) At both tolerances, multi-thread CPU parallelism via ma86 provides substantial speed-ups over the single-threaded solver Ipopt with ma27. (6) LiftedKKT and MadNCL.jl exhibit larger constraint violations compared to the other solvers.
@@
Figure 3: Proportion of MP OPF Cases Solved Within Factor of Fastest Solver
@@
No Storage $10^{-4}$ MadNLP (LK, GPU) 13 0.4 5.28e-4 9 7.09 2.59e-3 18 59.41 1.95e-3 40 17.33 1.63e-3
MadNLP (HK, GPU) 13 0.62 8.19e-7 8 18.09 7.34e-7 12 215.9 7.57e-7 33 42.32 7.76e-7
MadNCL (GPU) 13 0.63 2.36e-5 8 51.22 1.14e-2 14 225.28 3.89e-3 35 53.51 4.16e-3
Ipopt (CPU) 13 0.66 2.49e-6 7 132.77 8.41e-6 1 894.61 1.03e-6 21 131.0 4.40e-6
MadNLP (LK, CPU) 13 1.53 1.07e-6 8 86.38 4.93e-7 2 832.88 1.40e-5 23 118.27 1.99e-6
$10^{-8}$ MadNLP (LK, GPU) 13 0.95 5.28e-8 8 13.64 2.70e-7 12 239.19 1.56e-7 33 35.57 1.43e-7
MadNLP (HK, GPU) 13 0.52 4.99e-8 6 40.03 1.83e-7 8 259.11 1.86e-7 27 44.95 1.20e-7
MadNCL (GPU) 12 5.0 1.74e-7 5 89.39 4.15e-7 5 476.13 5.53e-7 22 82.91 3.15e-7
Ipopt (CPU) 13 0.89 4.95e-8 6 126.28 3.08e-7 0 - - 19 103.93 1.31e-7
MadNLP (LK, CPU) 13 3.39 5.28e-8 7 96.23 2.94e-7 1 833.42 4.19e-7 21 102.73 1.51e-7
Storage $10^{-4}$ MadNLP (LK, GPU) 16 3.13 5.70e-4 8 100.59 1.15e-3 15 305.46 8.94e-4 39 60.65 8.14e-4
MadNLP (HK, GPU) 10 49.8 2.35e-6 0 - - 0 - - 10 304.52 2.35e-6
MadNCL (GPU) 16 2.35 1.26e-4 9 137.53 8.66e-3 8 714.84 5.88e-4 33 90.32 2.56e-3
Ipopt (CPU) 15 12.14 1.83e-6 4 627.02 9.81e-8 0 - - 19 185.63 1.46e-6
MadNLP (LK, CPU) 15 14.64 8.27e-7 5 387.67 1.78e-7 0 - - 20 171.85 6.65e-7
$10^{-8}$ MadNLP (LK, GPU) 14 11.87 5.51e-8 6 119.75 1.53e-7 1 861.23 8.84e-8 21 38.34 8.47e-8
MadNLP (HK, GPU) 12 28.26 2.72e-8 0 - - 0 - - 12 109.73 2.72e-8
MadNCL (GPU) 14 10.57 5.44e-8 4 297.69 1.14e-7 1 821.55 6.39e-8 19 48.98 6.73e-8
Ipopt (CPU) 15 16.68 5.14e-8 1 642.34 6.23e-8 0 - - 16 76.59 5.21e-8
MadNLP (LK, CPU) 16 10.1 5.66e-8 4 446.25 8.86e-8 0 - - 20 55.37 6.30e-8
Cvio = maximum constraint violation, nvar = Number of Variables, LK = LiftedKKT, HK = HybridKKT.